### v1

# H-DES: a Quantum-Classical Hybrid Differential Equation Solver

Abstract
In this article, we introduce an original hybrid quantum-classical algorithm based on a variational quantum algorithm for solving systems of differential equations. The algorithm relies on a spectral method, which involves encoding the solution functions in the amplitudes of the quantum states generated by different parametrized circuits and transforms the task of solving the differential equations into an optimization problem. We first describe the principle of the algorithm from a theoretical point of view. We provide a detailed pseudo-code of the algorithm, on which we conduct a complexity analysis to highlight its scaling properties. We apply it to a set of examples, showcasing its applicability across diverse sets of differential equations. We discuss the advantages of our method and potential avenues for further exploration and refinement. [^1] The algorithm described in this paper is subject to a patent submission (EP24306601). Use of this method requires a license, which can be obtained from ColibriTD.
1 Introduction
The fine control of quantum systems is arguably among humanity’s most remarkable feats. The fragility of such systems, the core of the technological challenges faced by researchers, provides both an advantage and a disadvantage. The former motivates their use in quantum sensing beyond classical limits, as already demonstrated using gravitational wave detectors [YMT+20] and portable gravitometers with a sensitivity of over $10^{-9}g$ [MVL+18]. The latter hinders their use in simulations of quantum systems — a prospect first suggested by R. Feynman [Fey82]. Feynman’s idea was to use controlled quantum systems to simulate the behavior of natural systems. This task is commonly challenging for classical computers, depending on the system’s size and symmetries.
The simulation of Hamiltonians is the foundation of analog quantum computing [Zak99]. A relevant problem is mapped onto the dynamics of a quantum system, and the solution of the problem is usually encoded in the ground state of a Hamiltonian. The task of finding this ground state is particularly straightforward for quantum annealers, which apply the adiabatic theorem [BF28] to evolve an initial simple Hamiltonian into the final Hamiltonian. In fact, quantum annealers using the Ising model [Gal99] have already been manufactured and modest claims of quantum advantage have started to emerge [TAM+22]. This model can be translated into a quadratic unconstrained binary optimization (QUBO) problem [KHG+14], which finds numerous applications. QUBO problems have a simple formulation, while their optimal solutions, in contrast, are NP-hard to find by classical means. This approach can be adapted to many combinatorial optimization problems in finance [MKS+22], vehicle routing [BGK+20], logistics [SGM22], isomer search procedures [TANM20], and contact map overlaps [OSO18], to name a few. Other controlled quantum systems can be implemented using Rydberg atoms, which enable simulations of other dynamics, such as the XY Hamiltonian [HBS+20] and even ion traps, with more limited dynamics [ADG+22].
Controlled quantum systems find a more versatile (often universal, in a computational sense) use as quantum central processing units (QPUs). For this purpose, many algorithms have been developed that provide speedups over their best-performing classical counterparts, notably Shor’s factoring and Grover’s search algorithms [Sho97, Gro97]. Recent studies have begun to unravel the role played by superposition and entanglement in these speedups [JH19, BJH+20]. These algorithms, however, demand high qubit numbers and interconnectivity as well as long coherence times, much beyond what current noisy intermediate-scale quantum (NISQ) devices can offer. A promising alternative is provided by variational quantum algorithms (VQAs).
VQAs have the potential to perform well under NISQ-era limitations due to their moderate hardware requirements, among which are shallow parametrized circuits of $\lesssim 10^{2}$ qubits and the absence of all-to-all qubit gates [CAB+21]. The circuit parameters are optimized by a classical computer iteratively until the generated quantum state satisfies pre-imposed conditions(encoded in loss functions) [MBB+18]. Because of their potential for near-term applications, their integration with high-performance computing (HPC), and their similarity to machine learning techniques [WKS16], many such algorithms have been developed in several contexts. For example, the ground state of molecular Hamiltonians can be found using a variational quantum eigensolver (VQE) [TCC+21] by means of the Jordan–Wigner transformation; a system of linear equations can be solved using the variational quantum linear solver [BLC+19] as opposed to the NISQ-nonviable HHL algorithm [HHL09]; a variational version of the quantum phase estimation (QPE) algorithm has been developed in an effort to bring its usefulness into the NISQ era [FRF22]; and even combinatorial optimization problems are solvable via the quantum approximate optimization algorithm (QAOA) [FGG14].
Central to all these algorithms is the parametrized circuit structure, also called the Ansatz, as its choice may enable faster convergence. In VQEs, for instance, unitary coupled cluster Ansätze (plural of Ansatz) [PMS+14] are outperformed in some metrics by adaptive derivative-assembled pseudo-Trotter (ADAPT) ones [GEBM19], which use fewer gates, reducing classical optimization overheads. Nevertheless, the right choice of Ansatz, if it exists, may not be sufficient to ensure convergence, since the classical optimization may face barren plateaus, which prevent optimizers from finding the global minima of objective functions [WFC+21]. These objective functions are usually written in terms of expectation values, which need to be obtained efficiently to avoid shot noise. In these cases, the naive measurement of experimental outcomes can be replaced by approaches that require fewer shots, such as classical shadows [HKP20], its recent modifications [KG22], and Bayesian methods [MOB18, HH12, CG23].
Quantum algorithms have been developed not only for tackling problems in which classical computers perform poorly, but also for typically amenable problems, such as solving differential equations (DEs). To name a few: the heat equation was recently solved via a quantum annealer in a hybrid setup [PSG+21]; a fully quantum method was used to solve the wave equation [CJO19]; and another was used for the Navier–Stokes equations [Gai21]. Non-linear DEs such as the latter are among the most challenging to solve via numerical methods and have been the object of recently developed algorithms. In one instance, a non-linear Schrödinger equation was solved by a VQA algorithm that included a Hadamard test as a subroutine [LJM+20]. The latter approach is similar to that of classical finite difference methods (FDM) as the function’s domain is discretized. The authors claim $\sim 20$ error-corrected qubits ($2^{20}$ grid points) are enough to rival state-of-the-art supercomputers; the cost is optimizing a number of circuit parameters that grows linearly with the number of qubits. Another more versatile method resorts to differentiable quantum circuits (DQCs) [MNKF18, SBG+19] and quantum feature maps [MNKF18, LSI+20, SSM21] to evaluate derivatives without intrinsic numerical errors [KPE21]. One possible drawback is that the number of $n$-qubit circuits necessary grows as $\propto n^{m}$, where $m$ is the highest order of derivatives in the DE.
In this paper, we propose a variational quantum algorithm to solve partial DEs based on a spectral decomposition of the solution function. This approach avoids the problem of requiring an exponentially growing number of circuits to evaluate the derivative of the solution function but requires the efficient readout of expectation values. This article is organized as follows. In Sec. 2, we introduce the algorithm and its key principle and detail the steps of its workflow. In Sec. 3, we discuss the application of the algorithm to multivariate equations. In Sec. 4, we explain some theoretical elements concerning the complexity and scaling of the hybrid algorithm. In Sec. 5, we run the algorithm on several systems of DEs, showcasing the versatility of our method. In Sec. 6, we discuss the advantages of our algorithm and compare it to other approaches in the literature. In Sec. 7, we conclude the article and present possible improvements of our algorithm. In the appendices, we share how one can deal with boundary conditions in simple cases (see App. A), some examples to illustrate the theoretical ideas behind the algorithm (see App. B), more details from the complexity study (see App. C), and a theoretical framework for performance comparisons of several solvers (see App. D).
2 Description of the algorithm
In this section, we present the overall principle and detail the main steps of the algorithm. We explain the theoretical basis on which our algorithm is built, for which pedagogical examples can be found in App. B. For simplicity, we first discuss our algorithm in the context of ordinary DEs (ODEs) and then show in Sec. 3 how the method can naturally be extended to solve partial DEs (PDEs).
We start with an overview of the hybrid quantum-classical algorithm for solving DEs. We illustrate the general workflow of our algorithm in Fig. 1 and list the general steps below:
  1. 1.
    We take as input the PDEs and their boundary conditions;
  2. 2.
    We encode the problem in a loss function that gives us a measure of how far the current attempt is from the solution;
  3. 3.
    We generate trial states from parametrized VQCs;
  4. 4.
    We measure expectation values of observables with respect to these states to evaluate trial functions and their derivatives;
  5. 5.
    We compute the loss function using these evaluations;
  6. 6.
    Using a classical optimizer, we update the parameters of the circuits to minimize the loss function;
  7. 7.
    The hybrid loop, consisting of the four previous steps, is repeated until a given error tolerance is reached. The optimal parameters are then returned and the solution functions can be retrieved.
INPUT: PDEs VQC Expectation values Function evaluation Refer to caption Evaluate PDEs and BCs Loss function HYBRID LOOP OUTPUT: Optimal vqc Optimizeparameters
Figure 1: An illustration of the workflow of our approach.
We further detail the steps of the algorithm in the following subsections and provide a pseudo-code in Sec. 4.1.
2.1 Variational Quantum Circuit and Ansatz
To construct the parametrized state used to minimize the loss function, we need a variational quantum circuit (VQC). The architectures employed here are based on the structure of the hardware-efficient Ansatz (HEA). In the HEA, each set of parametrized single-qubit rotations is followed by an entangling layer of CNOT gates. This block of rotations and CNOTs is then repeated $d$ times, where $d$ is called the depth of the VQC. In what follows, we consider linear connectivity among qubits, but the CNOTs can be easily generalized to other connectivity architectures [LOCC24].
As $d$ increases, the number of rotation parameters $\theta$ increases proportionally, giving rise to a circuit with higher expressivity power. These layers make the VQC structure comparable to those of classical neural networks, with rotation gates playing the role of neurons and angles playing the role of synaptic weights. By tuning and adjusting the circuit’s parameters, we change the output state to minimize the loss function.
Generally, the rotation layer includes rotations around different axes, i.e., $R_{x}$, $R_{y}$, and $R_{z}$ gates, but in our case, this layer is composed only of parametrized $R_{y}$ gates for each qubit (see Fig. 2). This choice was driven by the fact that we do not need to generate states with complex amplitudes since only the associated probabilities are used to encode the solution function. We thus chose to restrict the research space to real states, for which using only $R_{y}$ gates is sufficient. We also empirically observed that results for different loss functions reached satisfactory precision with a single rotation gate per layer (at least in the ideal case, i.e., with perfect operations in the circuit).
[…]
Figure 2: A representation of the hardware-efficient Ansatz used in our variational quantum algorithm.
2.2 Spectral methods and Chebyshev polynomials
Our approach is based on the spectral method, in which the solution function is expressed as a linear combination of the elements of a complete basis of orthogonal functions. Therefore, the task of finding the best solution to a DE consists of determining the best approximation of the coefficients of the solution in that basis.
We choose to express the solution function in the basis of Chebyshev polynomials, which are widely used as basis functions for spectral methods [Boy01, WS07]. These polynomials form a complete orthogonal system and allow us to represent piecewise smooth and continuous functions in finite domains. The expression of their derivatives is straightforward and can also be defined by recurrence, making it easy and efficient to implement.
One can find different definitions of Chebyshev polynomials in the literature. In our algorithm, we use the “first kind”, defined as follows:
$\text{Cheb}(k,x)=\\ \begin{cases}\cos(k\arccos x)&\text{ if }|x|\leq 1\\ \cosh(k\operatorname{arcosh}x)&\text{ if }x\geq 1\\ (-1)^{k}\cosh(k\operatorname{arcosh}(-x))&\text{ if }x\leq-1\end{cases}~{},$ (1)
with $x\in\mathbb{R}$ the point at which we evaluate the polynomial and $k\in\mathbb{Z}^{+}$ the order of the Chebyshev polynomial.
Therefore, the expression of a given function $f:\mathbb{R}\to\mathbb{R}$ expressed in the truncated Chebyshev polynomial basis, with $C$ terms, can be written as follows:
$f(x)=\sum_{k=0}^{C-1}c_{k}~{}\text{Cheb}(k,x)~{},$ (2)
with $c_{0},c_{1},\dots,c_{C-1}\in\mathbb{R}$.
2.3 Encoding and evaluation of the function
In this subsection, we detail how we use the $n$-qubit state generated by the variational quantum circuit to represent the solution function. We denote as |$\ket{\psi_{f}}$ the state coming out of the VQC modeled by the gate $U_{\theta}$, expressed in the computational basis (with decimal notation) as follows:
|$\ket{\psi_{f}}=U_{\theta}\ket{0}^{\otimes n}=\sum_{i=0}^{2^{n}-1}a_{i}\ket{i}~{},$ (3)
with $\text{with }~{}~{}\sum_{i=0}^{2^{n}-1}|a_{i}|^{2}=1~{}.$ (4)
We encode the solution function $f$ in the amplitudes of the state, or more precisely, in the associated probabilities $p_{i}$. We associate a basis state with each Chebyshev polynomial (of a certain order), and the coefficient in front of this polynomial becomes the probability of measuring that basis state.
Since probabilities are positive real numbers, expressing $f$ using only positive coefficients in the basis would not allow us to represent all continuous functions. One would also need to combine positive and negative coefficients in the decomposition of the solution function. In order to do this, we add to the list of polynomial values a copy of each one, multiplied by $-1$. In the quantum encoding of our function, these probabilities are associated with the second half of the basis states (see Eq. 7).
Furthermore, in the particular case $x\in[-1,1]$, we have $\text{Cheb}(k,x)\in[-1,1]$ for all $x$ and $k$. When we take the linear combination of Chebyshev polynomials with the associated difference of probabilities (because of the duplication of Chebyshev polynomials with a minus sign), which are bounded by the normalization constrain, the resulting function is also bounded. Therefore, to allow for the modeling of any general solution function, we introduce a scaling parameter $\lambda\in\mathbb{R}$.
Finally, we encode the solution function $f$ in the following way:
$f(x)=\lambda\sum_{i=0}^{2^{n-1}-1}(p_{i}-p_{i+2^{n-1}})~{}\text{Cheb}(i,x)~{},$ (5)
with $p_{i}=|a_{i}|^{2}$.
Therefore, the variational circuit, for a given set of parameters $\theta=\{\theta_{i}\}_{i}$, models the state that encodes the decomposition of the function in the Chebyshev polynomial basis. Note that in the case of a system of DEs, each solution function has a different circuit and one can choose the number of qubits to allocate for each circuit independently.
To obtain or evaluate the solution function, one needs to retrieve the properties of its state. Since we are only concerned with the probabilities, there is no need to perform a full state tomography, since the probabilities correspond to the diagonal terms of the density matrix describing the state. Nevertheless, we still have to retrieve $N=2^{n}$ terms, which becomes difficult for $n\gg 1$ [MJZ+16, FB18, GXLK21, STN+14].
Instead, we determine the value of the solution $f$ at a point $x$, following the definition in Eq. 5, by computing the expectation value with respect to a diagonal observable $O_{C}(x)$ such that:
$𝜓𝑓f(x)=\lambda~{}\expectationvalue{O_{C}(x)}{\psi_{f}}~{},$ (6)
with
$\begin{gathered}O_{C}(x)=\sigma_{z}\otimes\left(\sum_{i=0}^{2^{n-1}-1}\text{Cheb}(i,x)\outerproduct{i}{i}\right)\\ =\begin{pmatrix}\text{Cheb}(0,x)\\ &\ddots&&&&\text{\huge 0}\\ &&\text{Cheb}(2^{n-1}-1,x)\\ &&&-\text{Cheb}(0,x)\\ \text{\huge 0}&&&&\ddots\\ &&&&&-\text{Cheb}(2^{n-1}-1,x)\\ \end{pmatrix}~{}.\end{gathered}$ (7)
By computing this expectation value, one can evaluate the function at a given point. This does not allow us to retrieve the full expression of the function $f$, but in the context of our VQA, we only need, at each iteration, to evaluate the function at specific points (see Sec. 2.5). We present a simple example to illustrate the encoding and evaluation of the functions in App. B.1.
2.4 Evaluation of the derivatives
One of the most remarkable advantages of this algorithm is the simplicity with which any derivative of a function can be calculated, and this is due to two characteristics of this method. The first is that a derivative of any order can expressed as
$\frac{\partial^{q}f(x)}{\partial x^{q}}=\lambda\sum_{i=0}^{2^{n-1}-1}(p_{i}-p_{i+2^{n-1}})~{}\frac{\partial^{q}\text{Cheb}(i,x)}{\partial x^{q}}$ (8)
and calculated by means of the same circuit as an expectation value of the observable
$𝜓𝑓\frac{\partial^{q}f(x)}{\partial x^{q}}=\lambda~{}\expectationvalue{O_{\partial^{q}C}(x)}{\psi_{f}}~{},$ (9)
where $O_{\partial^{q}C}(x)$ has exactly the same structure as $O_{C}(x)$ but with all the diagonal elements replaced by the $q$-order derivative of the corresponding Chebyshev polynomial. The second favorable characteristic is actually related to the Chebyshev polynomials directly, since any derivative ${\partial^{q}\text{Cheb}(i,x)}/{\partial x^{q}}$ can be defined using the following expression:
$\frac{\partial^{q}\text{Cheb}(i,x)}{\partial x^{q}}=2^{q}i\sideset{}{{}^{\prime}}{\sum}_{\begin{subarray}{c}0\leq k\leq i-q\\ k\equiv i-q~{}(\text{mod }2)\end{subarray}}\binom{\frac{i+q-k}{2}-1}{\frac{i-q-k}{2}}\dfrac{(\frac{i+q+k}{2}-1)!}{(\frac{i-q+k}{2})!}\text{Cheb}(k,x),$ (10)
where a prime after a summation symbol means that the term contributed by $k=0$ is to be halved, if it appears.
2.5 Loss function
The loss function $L$ encodes in itself the DEs we want to solve, giving us a way to quantify the difference between the solution of the problem considered and the trial functions we reconstruct by means of the observables. Using this value, the classical optimizer can update the parameters of the circuit with the ultimate aim of reducing this distance.
In order to build the loss-function, we collect all the terms of the DE on the left side, i.e., $E_{i}(x)=0$, where $E$ is the set of differential equation in our system, and we thus consider the $i^{th}$ differential equation. The evaluation of the DEs requires the evaluation of all the functions and their derivatives (appearing at least once) at each sample point, which in turn implies the computation of expectation values of the right observables.
Remark 1.
In the case of a linear DE involving only one function, one can evaluate it for a specific point $x_{s}$ by evaluating only one expectation value. In fact, the DE is defined as the linear combination of the solution function, its derivatives, and potentially the sample point itself. These terms are evaluated by computing the expectation value of an observable ($O_{\partial^{q}C}(x)$ for its $q^{th}$ derivative, and the scaled identity $x_{s}I$ for the sample point) on the same state. Therefore, the DE can be evaluated by regrouping all the observables into a single one using the linearity of the inner product.
The solution of a (system of) PDE(s) must respect both the functional and boundary conditions. These two criteria can be taken into account in the loss function as
$L(\theta)=L^{\rm{diff}}(\theta)+\eta\cdot L^{\rm{boundaries}}(\theta)~{},$ (11)
where the first term evaluates the match with respect to the DEs, while the second accounts for the boundary conditions.
The coefficient $\eta$ controls the weight of the boundary terms in the optimization and has to be chosen carefully to ensure that the constraint is imposed sufficiently. To tackle this task, called loss balancing in the context of multi-objective optimization, several strategies have been proposed in the deep-learning literature [BK21, HTM19, WTP21]. In some cases, it can be chosen so it is higher than the maximal possible values of the left term, and this represents a well-known problem in constrained optimization [WS07, NW06]. Adding a term directly to the loss function allows us to incorporate simple and complex boundary conditions within the same framework. When the boundary condition is simple or the functional form of the solution is reasonably independent of it, one can consider other ways of handling the boundary conditions (see App. A), and in that case we set $\eta=0$.
To construct both terms, we use the mean squared error (MSE). Using MSE metrics, the differential part of the loss function can be explicitly written as
$L^{\rm{diff}}(\theta)=\frac{1}{n_{s}}\sum_{e\in E}\sum_{x_{s}\in S}e(x_{s})^{2}~{},$ (12)
where the $x_{s}$ are the coordinates of the sample points, $\theta$ is the set of angles parametrizing the circuit. Similarly, the boundaries component is computed as the sum of all errors for each boundary condition for all functions appearing in the DE. This can be expressed as follows:
$L^{\rm{boundaries}}(\theta)=\frac{1}{n_{BC}}\sum_{f\in F}~{}\sum_{f_{BC},x_{BC}\in BC(f)}\big{(}f(x_{BC},\theta)-f_{BC}(x_{BC})\big{)}^{2}~{},$ (13)
with $x_{BC}$ and $f_{BC}$ the point and function, respectively, representing the boundary conditions for each function $f\in F$ and $n_{BC}$ the total number of boundary conditions.
2.6 Validation of the results
After retrieving the solution proposed by our hybrid solver, we want to evaluate the quality of the result. In this case, instead of replacing the solution functions in the DEs, as we do in the algorithm loop, we compare the solution to that provided by a classical solver (assuming that the latter returns the exact expression of the solution).
For each function $f\in F$ involved in the DEs to be solved, the expected solution $c_{f}$ is computed. The idea is to evaluate the quality of each solution function $f$ by computing some distance to $c_{f}$. We consider two definitions of the distance: $d_{1}$ and $d_{2}$. In both definitions, we compare the evaluation of the functions for a set of samples $S_{v}$, ideally containing many more elements than the set $S$ used for the optimization.
The first distance $d_{1}$ is computed by taking the maximum over $S_{v}$ of the absolute difference between the evaluation of the two functions:
$d_{1}(f,g)=\max_{x\in S_{v}}|f(x)-g(x)|~{}.$ (14)
The second distance is the average of the squared difference between the evaluation of the two functions throughout $S_{v}$. Using the average makes the comparisons of $d_{2}$s computed from differently sized $S_{v}$s more relevant.
$d_{2}(f,g)=\frac{1}{|S_{v}|}\sum_{x\in S_{v}}\big{(}f(x)-g(x)\big{)}^{2}~{}.$ (15)
Both distances provide different information about the quality of the fit. The first distance provides us with information about the maximum error of the solution function, which is useful if one needs a perfect fit at each sample point. The second distance indicates whether the solution overall is close to the true value, with potentially large or small variances in the errors. Therefore, by combining both distances, we can retrieve comprehensive information about the overall error as well as partial information on the variance. Let $V_{f,0}$ be the validation score for the function $f$, defined as the following tuple:
$V_{f,0}=\Big{(}d_{1}(f,c_{f}),~{}d_{2}(f,c_{f})\Big{)}~{}.$ (16)
However, fitting the function on the sample points does not necessarily ensure that the solution function, expressed in the spectral basis, satisfies the DE. One also has to ensure that the derivatives of the solution function (those appearing in the DE at least) match the derivatives of the actual solution. We thus extend the validation score $V_{f,0}$ to incorporate higher derivatives of the solution function as follows:
$V_{f,i}=\Big{(}d_{1}(f^{(i)},{c_{f}}^{(i)}),~{}d_{2}(f^{(i)},{c_{f}}^{(i)})\Big{)}~{},$ (17)
where $f^{(i)}$ is the $i^{th}$ derivative of the function $f$. One can then combine the validation scores of the function $f$ and its derivatives to define another overall score for the quality of the solution $f$. We denote as $G_{F}(f)$ the set of orders of derivatives of the function $f$ appearing in the set of DEs to solve and define the validation score $V_{f}$ such that:
$V_{f}=\Big{(}\max_{i\in G_{F}(f)}~{}V_{f,i}(0),~{}\frac{1}{|G_{F}(f)|}\sum_{i\in G_{F}(f)}V_{f,i}(1)\Big{)}~{},$ (18)
with $V_{f,i}(a)$ referring to the $a^{th}$ ($a\in\{0,1\}$) entry of the validation couple defined as $V_{f,i}$ in Eq. 17. This validation score $V_{f}$ gives us information on the quality of each solution function $f$ and can be compared with other validation scores since it is only defined using absolute differences, and their range of values is also independent of the number of samples in $S_{v}$.
We finally define a global validation score $V$ of the solver by combining each function’s validation score $V_{f}$, such that:
$V=\Big{(}\max_{f\in F}~{}V_{f}(0),~{}\frac{1}{|F|}\sum_{f\in F}V_{f}(1)\Big{)}~{}.$ (19)
The first value gives us information on the maximum error among all functions and sample points. Having a high first value warns us that for some points, the solution function can be very far from the expected result. The second value gives us information on the overall error among all functions and sample points. Having a low second value ensures that overall, the solution functions live around the true solution. If one value of the validation couple is not satisfying, one can compute the intermediate validation score $V_{f}$ to determine which solution function of the system of DEs is responsible for the error in the validation. If one wants to identify exactly whether the problem comes from a solution function or its derivatives, it is still possible to compute all the $V_{f,i}$ scores. Otherwise, if the global validation score satisfies some criteria or thresholds, it is sufficient to conclude that all the solutions and their derivatives are valid.
We thus obtain a validation score that can be used to compare solutions for different sets of DEs, different solver parameters, or different solvers entirely. The genericity of the validation score allows us to then establish an acceptance threshold below which we consider the set of DEs solved. This can be dictated by the precision needed in different fields or industries, or for instance by benchmarking several sets of DEs and tuning the threshold to match a given acceptance criteria. In App. D, we present a theoretical framework to describe and compare the quality of convergence of several solvers, within which the global validation score $V$ can be used to measure performance.
3 Multivariate functions
In this subsection, we detail the multivariate case, i.e., when a solution function $f$ takes as input $X=(x_{1},x_{2},\dots,x_{v})\in\mathbb{R}^{v}$. Since the variable dependence is encoded only in the observables, moving from 1-dimensional to $v$-dimensional space does not affect the rest of the algorithm.
In the uni-dimensional case, each basis state (of the computational basis) is used to represent, in binary notation, the order of a Chebyshev polynomial. Except for the first qubit, which is used to determine the sign in front of the polynomial, all the remaining qubits of the basis states are indeed used to encode the corresponding order. In the multivariate case, we assign to each variable $x_{j}$ a specific number of qubits that are used to encode the order of the corresponding Chebyshev polynomial. In fact, one possible way of defining the multivariate Chebyshev polynomial $\text{Cheb}(i,x_{1},x_{2},\dots,x_{v})$ is to take the product of independent univariate Chebyshev polynomials $\text{Cheb}(L_{j},x_{j})$ defined on a rectangular domain [PV11, Tre16, Mas80].
Therefore, each basis state |$\ket{i}$ of the $(n+1)$-qubit state modeling the solution function is used to encode the different indices assigned to each variable. We write the binary decomposition of the basis state |$\ket{i}$ (written in decimal notation) as
[…]
For each variable $x_{j}$, we attribute $l_{j}$ qubits to encode the order (of the Chebyshev polynomial) and denote as $L_{j}$ the binary decomposition of the order. This allows us to separate the basis state |$\ket{i}$ as the tensor product of the first (sign) qubit with $v$ states, such that
[…]
with
$l_{1}+l_{2}+\cdots+l_{v}=n~{}.$
In this way, we allow freedom in the selection of the required number of qubits $l_{j}$ for each variable $x_{j}$, which can be useful when some variables require more terms in the spectral decomposition. To summarize, we evaluate the multivariate Chebyshev polynomial using the following formula:
$\text{Cheb}(i,X)=\text{Cheb}(i,x_{1},x_{2},\dots,x_{v})=\prod_{j=1}^{v}\text{Cheb}(L_{j},x_{j})~{}.$
This allows us to write the corresponding encoding of a multivariate solution function $f$ with the same notation as before:
$f(x_{1},x_{2},\dots,x_{v})=\lambda\sum_{i=0}^{2^{n-1}-1}(p_{i}-p_{i+2^{n-1}})~{}\text{Cheb}(i,x_{1},x_{2},\dots,x_{v})~{}.$
Using the same idea, the only change appearing when evaluating the function at a sample point is in the expression of the observable $O_{C}$:
$𝜓𝑓f(x_{1},x_{2},\dots,x_{v})=\lambda~{}\expectationvalue{O_{C}(x_{1},x_{2},\dots,x_{v})}{\psi_{f}}~{},$
with
$O_{C}(x_{1},x_{2},\dots,x_{v})=\sigma_{z}\otimes\left(\sum_{i=0}^{2^{n-1}-1}\text{Cheb}(i,x_{1},x_{2},\dots,x_{v})\outerproduct{i}{i}\right)~{}.$
We present in App. B.3 a simple example to illustrate the encoding principle in the multivariate case.
3.1 Computing the derivatives efficiently
When we move to the multivariate case, the computation of the partial derivatives of the solution function can become more complex. However, since the multivariate Chebyshev polynomial is expressed as a product of independent single variable Chebyshev polynomials, it makes the derivatives easier to express. For instance, in the case of a partial derivative of order $g=1$, we can write:
$\frac{\partial\text{Cheb}}{\partial x_{j}}(i,x_{1},x_{2},\dots,x_{v})=\frac{\partial\text{Cheb}}{\partial x_{j}}(L_{j},x_{j})\times\prod_{k=1,k\neq j}^{v}\text{Cheb}(L_{k},x_{k})~{}.$ (21)
We point out that for computing the partial derivative with respect to the $j^{th}$ variable, we only need to compute one partial derivative. Furthermore, the rightmost term of the product corresponds to the product of all the remaining univariate Chebyshev polynomials that are already computed to evaluate the solution function. This large product is also equal to the value of the $v$-dimensional Chebyshev polynomial evaluated at the sample point of $\mathbb{R}^{v}$ divided by the univariate Chebyshev polynomial evaluated on the $j^{th}$ coordinate of the sample point. In other terms, we have
$\frac{\partial\text{Cheb}}{\partial x_{j}}(i,x_{1},x_{2},\dots,x_{v})=\frac{\partial\text{Cheb}}{\partial x_{j}}(L_{j},x_{j})\times\frac{\text{Cheb}(i,x_{1},x_{2},\dots,x_{v})}{\text{Cheb}(L_{j},x_{j})}~{},$ (22)
which allows to reduce the number of multiplications needed and reuse already computed evaluations of the Chebyshev polynomials (uni- and multi-variate). Therefore, by computing all first derivatives of the univariate Chebyshev polynomials, we can compute any first-order derivative of the multivariate Chebyshev polynomial.
When we move to the case of second-order derivatives ($g=2$), we need to distinguish between two cases: that of a partial derivative that is second order in a single variable and that of a partial derivative in two distinct variables. In the first case, we only need to elevate the derivative to second order as follows:
$\frac{\partial^{2}\text{Cheb}}{\partial{x_{j}}^{2}}(i,x_{1},x_{2},\dots,x_{v})=\frac{\partial^{2}\text{Cheb}}{\partial{x_{j}}^{2}}(L_{j},x_{j})\times\frac{\text{Cheb}(i,x_{1},x_{2},\dots,x_{v})}{\text{Cheb}(L_{j},x_{j})}~{}.$ (23)
Since the second case is a product of functions of independent variables, we retrieve the product of both derivatives with the product of the remaining Chebyshev polynomials. By recalling that $X=(x_{1},x_{2},\dots,x_{v})$, we obtain:
$\frac{\partial^{2}\text{Cheb}}{\partial x_{j_{1}}\partial x_{j_{2}}}(i,X)=\frac{\partial\text{Cheb}}{\partial x_{j_{1}}}(L_{j_{1}},x_{j_{1}})\times\frac{\partial\text{Cheb}}{\partial x_{j_{2}}}(L_{j_{2}},x_{j_{2}})\times\frac{\text{Cheb}(i,X)}{\text{Cheb}(L_{j_{1}},x_{j_{1}})\times\text{Cheb}(L_{j_{2}},x_{j_{2}})}~{}.$ (24)
This result can be generalized to any partial derivative of order $g$ with respect to the variables denoted by $x_{j_{1}},x_{j_{2}},\dots,x_{j_{g}}$. Some of these variables can appear several times in the partial derivative. We suppose that these $g$ variables can be regrouped into a set of $h$ $(\leq g)$ distinct variables $x_{k_{1}},x_{k_{2}},\dots,x_{k_{h}}$. We denote as $g_{l}$ the number of times the variable $x_{k_{l}}$ appears in the partial derivative, and we thus have $g_{1}+g_{2}+\cdots+g_{h}=g$. Therefore, we can rewrite the expression for the $g^{th}$ derivative in the following manner:
$\frac{\partial^{g}\text{Cheb}}{\partial x_{j_{1}}\dots\partial x_{j_{g}}}(i,X)=\frac{\partial^{g}\text{Cheb}}{\partial^{g_{1}}x_{k_{1}}\dots\partial^{g_{h}}x_{k_{h}}}(i,X)~{},$ (25)
such that:
$\frac{\partial^{g}\text{Cheb}}{\partial^{g_{1}}x_{k_{1}}\dots\partial^{g_{h}}x_{k_{h}}}(i,X)=\prod_{l=1}^{h}\frac{\partial^{g_{l}}\text{Cheb}}{\partial^{g_{l}}x_{k_{l}}}(L_{k_{l}},x_{k_{l}})\times\frac{\text{Cheb}(i,X)}{\prod_{l=1}^{h}\text{Cheb}(L_{k_{l}},x_{k_{l}})}~{}.$ (26)
In the case where $h$ is greater than half the number of variables, it is more efficient, for the second term of the product, to directly compute the product of the remaining variables instead of dividing the multivariate Chebyshev polynomial, which leads us to:
$\frac{\partial^{g}\text{Cheb}}{\partial^{g_{1}}x_{k_{1}}\dots\partial^{g_{h}}x_{k_{h}}}(i,X)=\prod_{l=1}^{h}\frac{\partial^{g_{l}}\text{Cheb}}{\partial^{g_{l}}x_{k_{l}}}(L_{k_{l}},x_{k_{l}})\times\prod_{j\notin\{k_{l}\}_{l}}\text{Cheb}(L_{j},x_{j})~{}.$ (27)
[…]
In the worst case, we can imagine that all the possible partial derivatives are involved in the DEs and that we will have to compute them all. In this scenario (supposing that we already computed all the Chebyshev functions needed), we must compute
$\sum_{g=1}^{n_{d}}\binom{n_{v}+g-1}{n_{v}}$ (28)
partial derivatives, which correspond to the sum of all possible partial derivatives at each order, with $n_{d}=\max(G_{F})$ the highest derivative order involved in the DEs and $n_{v}$ the maximum number of variables in functions involved in the DEs (see next section for the definition of $G_{F}$). Consequently, the computational cost of computing all the partial derivatives, expressed as a function of the cost $C_{b}$ of a multiplication/division, is equal to:
$\sum_{g=1}^{n_{d}}\binom{n_{v}+g-1}{n_{v}}\times(2h_{d}+1)\times C_{b}=\frac{n_{d}\times\binom{n_{d}+n_{v}}{n_{v}}}{n_{v}+1}\times(2h_{d}+1)\times C_{b}~{}.$ (29)
To compute all these partial derivatives, we only need to evaluate and multiply the value of the $d^{th}$ partial derivatives in one variable $\frac{\partial^{d}\text{Cheb}}{\partial{x_{j}}^{d}}(L_{j},x_{j})$. For each value of $d$, we have $n_{v}$ such partial derivatives to compute. In the worst case, we have to prepare all the possible partial derivatives of that form. This means that, for half of the basis states (since the other half only introduce a minus sign in front of the Chebyshev polynomials) and all sample points, in the worst case, we compute and store $2^{n-1}\times n_{v}\times n_{d}\times n_{s}$ values.
4 Implementation and related complexity
In this section, we conduct a theoretical study of our algorithm to examine the complexity of each step. More precisely, we aim to use theoretical arguments to study the algorithm scaling properties (for both the quantum and classical parts) as a function of the inputs of the problem to be solved (e.g., the number of sample points, number of variables, and number of DEs). This analysis will allow us to ensure the stability and realism of our method when moving to more complex systems of DEs. In the following, we will mainly focus on establishing formulas to estimate the computational time.
4.1 Pseudo-code of the algorithm
In this subsection, we present a pseudo-code (1) of the algorithm listing its main steps. The algorithm takes as input all information needed to define the differential equations to solve, as well as the required number of qubits $n$, the depth of the circuit $d$, the set of sample points $S$, the number of sample points $|S|=n_{s}$, and the target precision $\varepsilon$ for the loss function. The set $F$ contains the set of all functions that appear in the set of differential equations $E$. We denote as $G_{F}$ the set of all orders of derivatives (using an order of 0 for the function itself) appearing in $E$ for each function in $F$. We denote as p the parameters of the VQC that are inputs of the EVAL function.
Algorithm 1 H-DES (Hybrid Differential Equation Solver)
1:$E$ : the set of differential equations to solve
2:$F$ : the set of functions involved in $E$
3:$G_{F}$ : the set of orders of derivatives involved in $E$ for each function in $F$
4:$D$ : the set of domains of definition for each function in $F$
5:$n$ : the number of qubits
6:$d$ : the depth of the VQC
7:$n_{s}$ : the number of sample points
8:$\varepsilon$ : the target precision for the error
9:$m_{iter}$ : the maximum number of iterations of the optimizer
10:Approximation of every function $f\in F$
11:listObs, S $\leftarrow$ GenerateObservables($G_{F},D,n,n_{s}$)
12:qc $\leftarrow$ VQC($n$, $d$)
13:function Eval(p)
[…]
15:     return loss_function($E$, $S$, evals)
16:end function
17:p0 $\leftarrow$ initParams($E,F,D,n,d$)
18:p $\leftarrow$ optimizer(Eval, p0)
19:solF $\leftarrow$ reconstructFunctions($n$, $d$, p)
20:return solF
The algorithm returns the solutions of the differential equations in an approximate analytical form, which corresponds to the decomposition of the functions in the Chebychev basis.
4.2 Line-by-line analysis
In this section, we detail, line-by-line, the steps of Algorithm 1 and the related cost. For more details on some steps, we refer the reader to the appropriate section in Appendix C.
  • In line 11, the first classical pre-processing step is to build all the observables that will be used later in the algorithm to evaluate the functions and their derivatives, which requires the evaluation of Chebyshev polynomials at specific sample points. The analysis of this function is treated in Appendix C.1, and the time for completion is $T_{obs}=|G|\times n_{s}\times(2^{n-1}\times T_{cheb}+T_{diag})$, where $G$ is defined in the mentioned appendix.
  • In line 12, we build the parametrized quantum circuits. This step is detailed in Appendix C.2 and takes a time $T_{VQC\_init}=(2n-1)\times d\times(T_{gate\_creation}+T_{instruction\_adding})$.
  • In line 14, we compute the expectation value of the circuit for each observable. This computation is usually delegated to the quantum devices that decompose the observable and select the measurements to optimize this process. We discuss this in more detail in Appendix C.3, and this step takes a time $T_{expvalue}=n_{shots}\times(T_{VQC\_exec}+T_{measurement})+T_{post\_estim}$.
  • In line 15, we compute the loss function. This step is detailed in Appendix C.4 and takes a time $T_{loss}=|E|\times n_{s}\times(T_{square}+\sum_{f\in F}|H_{E}|T_{sum})$.
  • In line 17, we initialize the angles and scaling factor for all functions using a uniform random number generator. The associated cost is $T_{param\_init}=n_{f}\times(n\times d+1)\times T_{rand}$.
  • In line 18, we run a classical optimizer on the Eval function. For BFGS, it has in general a complexity of $\mathcal{O}(k^{2})$ [NW06], with $k$ the number of parameters to optimize, which in our case gives us $\mathcal{O}({n_{f}}^{2}\times n^{2}\times d^{2})$.
  • In line 19, we retrieve the solution functions by running the VQC with the optimal parameters on a simulator and retrieving the probabilities of the quantum state, which give us the decomposition of the function in the Chebychev basis as an approximate analytical form. We consider this a classical post-processing step that does not affect the efficiency of the optimization algorithm.
4.3 Scaling properties of the algorithm
In this subsection, we detail how the resources and parameters of the algorithm scale when we increase one specific parameter or move to more complex DEs.
First, when the number of variables increases, for instance when moving from ODEs to PDEs, only the pre-processing part, the number of qubits, and the number of samples are potentially affected. Indeed, we may need to increase the number of sample points in the research space to also include variations in the additional introduced variables. In the pre-processing part, it also implies the evaluation of additional univariate Chebyshev polynomials as well as additional multiplications (in the generation of the diagonal terms of the observable). Furthermore, if we want to maintain a specific precision for each variable of a function, we must multiply the number of qubits in the circuit by the number of variables. In other words, if one needs $m$ qubits for each variable, the number of qubits should satisfy $n=n_{v}\times m$, which clearly grows linearly with the number of variables. If one decides that the number of qubits must be a fixed value, then when the number of variables increases, the qubits (and thus the precision) for each variable are distributed equally or not, depending on the divisibility of $n$.
When the number of DEs increases, the number of solution functions naturally increases, leading to a proportional increase in various components of the algorithm. In fact, when we add functions, we need to add the corresponding number of variational quantum circuits, which also introduces additional parameters to optimize (additional angles and scaling factors). This can affect the performance of the (BFGS) optimizer quadratically. Having more circuits to evaluate also requires the evaluation of the expectation values (or at least the Pauli observables that decompose the diagonal observable) for these new circuits. However, the execution of the different circuits can be performed in parallel in a single iteration of the hybrid loop, since the parameters of the different circuits are independent. Finally, increasing the number of equations to solve implies a proportional increase in summing the terms of the loss function (and some additional boundary conditions to handle).
An increase in the order of the DE mainly affects the observables. First, it can require the evaluation of additional derivatives of the Chebyshev polynomials (if we suppose that all lower orders of derivatives appear in the DEs) in the classical pre-processing step, but the increase is proportional to the number of samples $n_{s}$. In addition, it can require the evaluation of more expectation values but this only affects the post-processing of such expectation values at each iteration, because they can also be deduced from the expectation value of the Pauli observables that are combinations of $I$ and $Z$, and thus do not require additional circuit measurements. The same argument for expectation values can be used when we increase the number of terms in the DEs. However, in this latter case, the evaluation of the equations for computing the loss consequently increases (also linearly).
Moving from a linear to a non-linear DE does not substantially change either the principle or complexity of the algorithm, and this is a notable strength of our approach. However, as mentioned in Rem. 1, one can no longer regroup the expectation values of the different observables into a single one, but this only affects the post-processing of the measurements after each iteration and depends on the number of terms (derivatives, functions, variables, constants) in the DE.
When we increase the number of qubits, the number of Chebyshev polynomials in the decomposition of the solution function increases exponentially, leading to potentially better solution precision. Furthermore, the number of parameters to optimize increases linearly with the number of qubits, which can affect the performance of the (BFGS) optimizer quadratically. It also affects the variational circuit, since for each additional qubit, we add a single-qubit rotation gate and another CNOT gate for each layer. However, this approach is not designed to require a large number of qubits, since with six (or seven) qubits, we already retrieve a spectral basis of size 32 (or 64), which is sufficient to express a large set of continuous functions.
When the depth of the circuit increases, more gates are executed, which linearly increases both the execution time and the number of parameters to optimize, quadratically affecting the performance of the optimizer. The idea behind a VQA is to work with shallow circuits and in our context of low-qubit circuits, choosing an appropriate Ansatz can ensure that the required depth for good expressivity is also low [DTYT22].
An increase in the number of sample points requires the manipulation of more observables to evaluate the function and its derivatives at those points. Evaluating the Chebyshev polynomials at each sample point linearly increases the pre-processing time, as well as the post-processing of the expectation values after each iteration of the algorithm. Finally, it also increases the complexity of computing the loss function proportionally.
The final consideration is what happens when we want to increase the precision required for the final loss function. The precision depends on several parameters mentioned above, and to achieve better precision, one would need to increase the number of iterations of the optimizer, increase the number of qubits, or increase the number of sample points, depending on where the algorithm needs help.
5 Results
In this section, we present the results obtained when applying our algorithm to solve different ODEs. We consider three examples: the first is a system of two coupled linear and first-order ODEs; the second is the damped harmonic oscillator, represented by a linear second-order ODE; and the third is a hypoelastic deformation problem modeled by two coupled non-linear and first-order ODEs.
We run our algorithm on the ideal QLM simulator ($qlm\_35$) provided by Eviden and give the size and depth of the circuit for each example. To generate a final state solution, we use the BFGS [NW06]*Sec. 6.1 optimizer to update the circuit angles $\theta$ within our hybrid loop. The following results are obtained by averaging over 100 random(uniform) initializations of the parameters of the circuits. We present the results in plots showing the solution function at each sample point as computed by both our solver and a classical (Mathematica DSolve function) solver. For each point, we also compute and plot the standard deviation. Finally, we present the validation scores associated with each solution, as defined in Sec. 2.6. We compute the validation score using a set $S_{v}$ of 100 points, linearly spaced in the interval $[0,0.95]$.
5.1 System of two coupled linear first-order ODEs
The first example we solve here is the following system of two linear coupled DEs:
$\begin{split}\frac{df(x)}{dx}-5=0~{},\\ \frac{dg(x)}{dx}-f(x)-5=0~{},\end{split}$ (30)
with the boundary conditions $f(0)=0$ and $g(0)=0$. To get valuable results, we estimated that 4 qubits, an Ansatz of depth 3, and 150 iterations with the BFGS optimizer are sufficient. In Fig. 3, we compare the attempted solution with the analytical results. We obtain a global validation score of $V=(1.95\cdot 10^{-3},6.20\cdot 10^{-7})$ and a final loss equal to $5.12\cdot 10^{-5}$.
Figure 3: The exact (blue solid) versus attempted solutions (orange dashed) for $f(x)$ (left) and $g(x)$ (right) mediated over 100 attempts after 150 iterations of the BFGS optimizer. The fit was obtained by minimizing the loss function over 20 points equally spaced in the $x=[0,0.95]$ domain. The area at the bottom shows the standard deviation (between the attempted and target solution) associated with each point.
5.2 Damped harmonic oscillator
Figure 4: A comparison between the target function (blue solid) obtained analytically and the attempted function of the algorithm (orange dashed). The area at the bottom shows the standard deviation (between the two functions) associated with each point.
The equation describing the damping of a harmonic oscillator is a linear second-order ODE:
$\frac{d^{2}x}{dt}+2\zeta\omega\frac{dx}{dt}+\omega^{2}x=0~{},$ (31)
with the boundary conditions $x(0)=2$ and $x^{\prime}(0)=0$. Here, $\omega$ represents the undamped angular frequency of the oscillator and $\zeta$ the damping ratio, and these values are respectively set to $\frac{9}{8}$ and $\frac{45}{8}$. We estimate that, to get valuable results, five qubits, an Ansatz of depth 5, and 525 iterations with a BFGS optimizer are needed.
In Fig. 4, we compare the attempted solution with the analytical results. We obtain a global validation score of $V=(2.87\cdot 10^{-2},3.88\cdot 10^{-4})$ and a final loss equal to $2.69\cdot 10^{-3}$.
5.3 Material deformation application
The last example we consider to demonstrate the performance of our algorithm is a material deformation problem. We consider simple and coupled DEs to describe one-dimensional samples, such as a tensile test, where a thin strip is fixed to a grip and pulled at the right end. One of the goals of solving DEs for static material deformation is to compute the final displacement of the solid under a load. This involves determining the spatial difference for each point between the initial and deformed solids.
Generally, material deformation can be described in terms of forces per unit area applied to the sample, called stress, and the resulting strain, which characterizes the relative displacement of the body with respect to the reference length. Mathematically, this is represented by relating a stress tensor $\sigma_{ab}$ and a strain tensor $\epsilon_{ab}$, where $a,b=x,y,z$ denote the directions in Cartesian coordinates. This gives a set of DEs whose form (ordinary or partial, linear or non-linear, etc.) depends on the problem considered and approximations applied.
As a proof of principle, we consider a one-dimensional case of material deformation in the hypoelastic regime. Hypoelastic refers to a non-linear, yet reversible, stress-strain constitutive relation. This means the deformations (described by the strain $\epsilon_{xx}$) respond non-linearly to external forces (the stress $\sigma_{xx}$). The governing non-linear coupled ordinary DEs are:
[…]
where $u(x)$ is the displacement along $x$ and $b_{x}$ is the constant change in the stress along the $x$-direction. The stress-strain constitutive relation is given by
$\epsilon_{xx}=\frac{\sigma_{xx}}{3K}+\frac{2\epsilon_{0}}{\sqrt{3}}\left(\frac{\sigma_{xx}}{\sqrt{3}\sigma_{0}}\right)^{n},$ (33)
where $b$, $K$, $\sigma_{0}$, and $\epsilon_{0}$ are material parameters defined by the problem. We consider a thin strip of length $L$ fixed to a grip at the left end and pulled at the right end. The corresponding boundary conditions are $u(0)\equiv u_{0}=0$ and $\sigma_{xx}(L)\equiv t$.
To simulate tensile testing (a fundamental test in material science and engineering) of a metallic strip, we set the material parameters to $b=10$, $\sigma_{0}=5$, $\epsilon_{0}=0.1$, $n=4$, and $K=100$, and we use $L=0.9$ and $t=2$ with 20 equally spaced points to sample the strip. We run our algorithm with 4 qubits, an Ansatz of depth 3, and 400 iterations of the optimizer. In Fig. 5, we compare the attempted solution with the analytical solutions. We obtain a global validation score of $V=(2.59\cdot 10^{-2},3.34\cdot 10^{-4})$ and a final loss equal to $1.05\cdot 10^{-3}$.
Figure 5: The exact (blue solid) versus attempted solutions (orange dashed) for the displacement (left) and stress (right) average over 100 attempts and after 400 iterations of the BFGS optimizer. The fit was obtained by minimizing the loss function over 20 points equally spaced in the $x=[0,0.95]$ domain. The area at the bottom shows the standard deviation associated with each point.
6 Discussion
6.1 Advantages of our approach
The first advantage of our method is that the number of qubits required is limited. First, the representativity power of our approach grows exponentially with the number of qubits, since the number of Chebyshev polynomials doubles when we add one qubit. With an already low number of qubits, one can represent a large range of continuous functions. Second, the qubit requirement grows linearly with the number of dimensions. The latter is a crucial point: since the number of variables is usually limited by spatial and temporal dimensions (3D$+$1D), the number of qubits needed to represent the solution function is also limited.
Moreover, the number of circuits needed in this approach also remains small. In fact, only one circuit for each iteration of the loop is required to evaluate both a function and its derivatives. This means that the number of circuits needed is exactly equal to the number of functions involved in a system of partial DEs.
Compared to a usual polynomial interpolation in a basis of size $2^{n}$, we have exponentially fewer parameters to optimize due to the exponential dimension of the $n$-qubit Hilbert space. Indeed, with the variational quantum circuit of $n$ qubits presented before, we only have to optimize $n\times d$ angles (with $d$ the depth). Therefore, the number of parameters to optimize increases linearly with the number of qubits.
In addition, as described before, good results using shallow circuits ($d=3$ for 1D MD) have been obtained. The expressivity provided by the parametrized quantum circuit allows us to find the solution without a need for deep circuits.
Finally, another advantage of our method is the lack of error in the computation of the derivatives. Since the same circuit is used to define the state representing the function and its derivatives, no approximation is needed to estimate them. For these reasons, we believe that the main expected advantage provided by the algorithm is its potential to return very precise solutions with a limited amount of quantum resources. It is also possible that this advantage translates into a reduction in energy consumption for a fixed precision when compared to classical solvers.
6.2 Differentiation from existing algorithms
Many quantum algorithms have been proposed in the literature for solving DEs. However, most of the proposed methods suffer from limitations that do not allow them to be good candidates for a general quantum DE solver. In this section, we list the state-of-the-art quantum algorithms for solving DEs and regroup them in Tab. 1, according to their main principles and limitations. Most of the quantum algorithms presented are restricted in terms of types of DEs that they can solve, and the majority of the listed methods rely on the finite difference method which introduce approximations in the derivatives of the trial functions. Our method has the advantage to tackle any system of partial DEs in the NISQ era, while using an exact expression of the derivatives.
Specificity References
One restricted type of equation (Burger, Vlasov, …) [GJFC, ESP19, OVKG22]
Solves only linear ODEs [Ber14, BCOW17, XWC+20, WWL+, CL20]
Solves only ODEs [SV21, LO08, Kac06, ZMSS21]
Solves only linear PDEs [FJO21, FML17, SS19, CLO21, KDE+20, AKWL19, CJO19, LMS20, FL18, OVKG22]
Solves only polynomial DEs [LO08]
Transforms DEs into systems of linear equations [SS19, BNWA23a, CLO21, PSG+21, KDE+20, LDG+20, CJO19, Bud21, MP16]
(using FEM, FDM, QFT) [ZMSS21, WWL+, GRG22, CL20, FL18, OVKG22, LJM+20, CLO21, LJM+20]
Transforms PDE into ODEs [Gai21, MK22b]
(Semi-)spectral approach [CL20]
Not NISQ friendly [Gai21, Ber14, CLO21, Bud21, Ste20, GRG22]
Other paradigm (annealing) [SS19, PSG+21]
Table 1: Specificities of state-of-the-art quantum algorithms for solving DEs. FEM stands for Finite Elements Method, while QFT stands for Quantum Fourier Transform.
We can compare our work in particular to that of Kyriienko et al. [KPE21], who proposed a general ODE solver based on differentiable quantum circuits. In our approach, we encode the dependence on $x$ in the observables, while Kyriienko et al. encode it directly in the circuit using quantum feature maps. In their approach, the evaluation of the solution function is a non-linear process, allowing different choices of observables of quantum feature maps, but not based on a specific encoding or decomposition of the solution (i.e., it is not a spectral approach). For evaluating the solution function or its derivatives, we replace Chebyshev polynomials with their derivatives in the diagonal observables, while the approach of Kyriienko et al. involves reconstructing the derivatives by evaluating several circuits (changing the quantum feature map of the circuit used to evaluate the solution function), whose number grows exponentially with the order of the derivative. However, Kyriienko et al. do not discuss in detail how to solve systems of DEs, or partial DEs and the encoding of multi-dimensional variables, for which the cost and its scaling are also not precised.
7 Conclusion
In this article, we have proposed an original hybrid quantum algorithm for solving partial DEs. By encoding the solution functions in parametrized quantum circuits and leveraging the principle of variational quantum algorithms, we have transformed the problem of solving such equations into an optimization problem. Using a spectral method, the algorithm is able to evaluate the solution function and its derivatives from a single parametrized circuit by computing expectation values of variable-dependent observables without approximating the derivatives needed. The solver optimizes a loss function that measures how the DEs are fit on a set of sample points.
The algorithm is designed to handle any linear or non-linear DE. As detailed in previous sections, its ability to be easily implemented in the case of partial and coupled DEs demonstrates the genericity of the approach and the versatility of our solver. Its scalability properties show that adding more complexity to the problem by introducing non-linear terms, more variables, or more functions to the system of DEs only linearly increases the number of resources (qubits or circuits) and parameters to optimize. In addition, with the expressivity of the solution function increasing exponentially with the number of qubits, and with shallow circuits needed to encode it, the algorithm is suitable for running on current quantum devices. We thus believe this algorithm constitutes an important step toward toward realistic applications of hybrid quantum algorithms
The natural next steps will be to demonstrate the ability of our algorithm to perform in a noisy environment and to develop a full pipeline including error mitigation techniques. Another possible improvement of the algorithm will be to provide an initialization strategy for the circuit parameters, for instance by exploiting the expression or properties of the DEs to start the optimization process as close as possible to the actual solution. In the direction of having a generic solver, one can imagine that some parameters of the solver (the number of qubits, depth, number of iterations, number of sample points) can be automatically chosen to satisfy the desired precision by taking into account the type and form of the DE. Improvements to the algorithm can be made at different stages to optimize some aspects of pre and post processing, to more efficiently evaluate the expectation values of the diagonal observables, and to provide a loss function less prone to barren plateaus and local minima, but also by selecting an appropriate classical optimizer.
In conclusion, the hybrid quantum algorithm proposed in this paper exhibits promising potential for practical application to various use cases despite the current limitations inherent to quantum computing. By combining classical computational methods with nascent quantum capabilities, this algorithm provides an original solution for solving generic partial DEs. In the NISQ era, where quantum computers face challenges related to stability and computational power, the significance of this hybrid approach relies on the potential span of diverse sectors, from material science to finance, where its deployment could help optimize processes such as material design and risk assessment. This hybrid algorithm not only serves as a bridge between theoretical quantum advancements and industrial applicability but also underscores the current need for accelerated integration of quantum technologies into industrial workflows. We believe that the adoption of such algorithms will signify a critical stride toward realizing the transformative potential of quantum computation in practical industrial contexts.
Acknowledgments
The authors want to thank Gözde Üstün, Xavier Pillet, and Ambroise Müller for constructive discussions on a preliminary version of the algorithm. The authors also thank Laurent Guiraud, Youcef Mohdeb, Muhammad Attallah, Roland Katz, Hugo Bartolomei, Roman Randrianarisoa and Aoife Boyle for their valuable insights and feedback, which significantly enhanced the quality of this article.
Appendix A Boundary conditions within the sample points
In order to deal with simple boundary conditions, we first need to distinguish between two cases: those in which the boundary condition concerns the function $f(x)$ itself and those in which it is associated with a $q$-order derivative of the function. In the first case, we use a method called floating boundary conditions handling, while in the second case, exactly like for finite difference methods, we exploit the discretization to manipulate $q+1$ points around the coordinate $x_{0}$ of the boundary condition in order to make them match the derivative at a certain point. The details of these methods will be presented in the following.
A.1 Floating boundary handling
This method consists of iteratively shifting the attempt solution to force matching of the boundary conditions. Including boundary conditions through this method allows us to solve the DE shifted to any position, consequently enlarging the set of possible solutions. For the sake of generality, we consider the case in which $n_{BC}$ boundary conditions at points $BC(f)=\{x_{j}\}_{j}$ need to be applied to a function $f$, i.e.,
$\begin{split}f(x_{1})&=k_{1}~{},\\ f(x_{2})&=k_{2}~{},\\ &\vdots\\ f(x_{n_{BC}})&=k_{n_{BC}}~{}.\end{split}$ (34)
Let us denote as $\tilde{f}$ the attempt solution of the current iteration. We associate each point in $BC(f)$ with a shift value as follows:
$y_{j}=\tilde{f}(x_{j})-k_{j}$ (35)
We then perform a polynomial fit on the points $\{(x_{j},y_{j})\}_{j}$ of order $n_{BC}-1$, resulting in a function shift, such that $\text{shift}(x_{j})=y_{j}$ for $x_{j}\in BC(f)$. The new shifted attempt function is expressed as
$\tilde{f}_{\text{shifted}}(x)=\tilde{f}(x)-\text{shift}(x)~{}.$ (36)
Furthermore, before evaluating the loss function at each iteration, we also have to coherently adjust each $q$-order derivative by adding the corresponding term $\partial^{q}\text{shift}/\partial x^{q}$. A concrete example can be found in App. B.2.
The main difference between this method and that including the boundary conditions directly within the loss function is that the first takes the boundary conditions into account exactly and with certitude, since the value of the solution function is manually fixed. The other solution can only perfectly ensure the boundary condition is met if the loss function is exactly equal to zero. Having a slight imperfection in the boundary condition fit can lead to very different solutions, especially in a chaotic setup. However, floating boundary conditions handling cannot be used when the boundary conditions are more complex.
A.2 Tangential approximation
In order to impose boundary conditions on certain derivatives, we can also consider a method called the tangential approximation or the finite difference trick. The idea is to use the finite difference approximation of the derivatives to impose values of the solution function on a set of close points. For example, if a DE has a boundary condition on the first derivative of $f$ at $x_{0}$ of value $k_{0}$, one can write the following approximation:
$\frac{\partial f}{\partial x}(x_{0})=k_{0}\approx\frac{f(x_{1})-f(x_{0})}{x_{1}-x_{0}}~{},$ (37)
with $x_{1}$ a point chosen to be close enough to $x_{0}$. If we denote the attempt solution at the current iteration as $\tilde{f}$, one can impose the value of the function at $x_{0}$ such that:
$\tilde{f}(x_{0})=\tilde{f}(x_{1})-k_{0}(x_{1}-x_{0})~{}.$ (38)
By adding a point $x_{1}$ to the sample points $S$ of the algorithm and also including the point $x_{0}$ in the evaluation of the loss function(and forcing the value according to Eq. 38), we transform the problem of satisfying the boundary condition in $\frac{\partial f}{\partial x}(x_{0})$ into optimizing $\tilde{f}(x_{1})$ such that the finite difference is as close as possible to the value $k_{0}$.
Once this is done, similarly to the shift presented in the previous section, we manually replace the value of the derivative $\partial f/\partial x$ before computing the loss function value of the attempt function $\tilde{f}(x)$. It is worth noting that if the boundary applies to the point $x_{0}$, we use backward derivative discretization; otherwise, we use forward derivative discretization. Unlike floating boundary handling, this method is not exact, and its efficiency increases as we increase the number of sample points or at least reduce the distance between the points involved in the process.
Appendix B Examples
B.1 Modeling a given function
In this section, we provide simple examples to illustrate the encoding of the solution function in the quantum state generated by the VQC and the evaluation of the function using observables depending on Chebyshev polynomials.
Let $f:\mathbb{R}\to\mathbb{R}$ be the function defined by $f(x)=2x-3$. Since $f$ is a linear polynomial in $x$ of degree 1, it can be exactly expressed using the two first Chebyshev polynomials, i.e.,
$\text{Cheb}(0,x)=\cos(0)=1~{},~{}~{}~{}~{}~{}\\ \text{Cheb}(1,x)=x~{}.$ (39)
Therefore, we can express $f$ as
$f(x)=-3\cdot\text{Cheb}(0,x)+2\cdot\text{Cheb}(1,x)~{},$ (40)
which we can rewrite in the form of Eq. 5 as follows:
$f(x)=5\left(\Big{(}0-\frac{3}{5}\Big{)}\text{Cheb}(0,x)+\Big{(}\frac{2}{5}-0\Big{)}\cdot\text{Cheb}(1,x)\right)~{}.$ (41)
We see clearly from this expression that $f$ can be modeled using the global scaling factor $\lambda_{f}$ and the two-qubit quantum state |$\ket{\psi_{f}}$ such that:
$\lambda_{f}$ $=5$ (42)
|$\ket{\psi_{f}}$ $=\frac{1}{\sqrt{5}}\left(\sqrt{2}\ket{01}+\sqrt{3}\ket{10}\right).$ (43)
One can thus write
$f(x)=\lambda_{f}\langle\psi_{f}|O_{C}(x)|\psi_{f}\rangle$ (44)
with
$O_{C}(x)=\begin{pmatrix}1&0&0&0\\ 0&x&0&0\\ 0&0&-1&0\\ 0&0&0&-x\end{pmatrix}~{}.$ (45)
B.2 Floating boundary handling
In this section, we give a concrete example of how to handle boundary conditions on a certain function $f:\mathbb{R}\to\mathbb{R}$. Let us suppose we have two boundary conditions:
$\begin{split}f(x_{0})&=a~{},\\ f(x_{1})&=b~{}.\end{split}$ (46)
At each step of the optimization process, we shift the attempt solution $\tilde{f}(x)$ by constructing a straight line shift function using the following values:
$\begin{split}\text{shift}(x_{0})&=\tilde{f}(x_{0})-f(x_{0})=\tilde{f}(x_{0})-a~{},\\ \text{shift}(x_{1})&=\tilde{f}(x_{1})-f(x_{1})=\tilde{f}(x_{0})-b~{}.\end{split}$ (47)
Then, the new attempt function is:
$\tilde{f}_{\text{shift}}(x)=\tilde{f}(x)-\text{shift}(x).$ (48)
In this specific example, since we are dealing with a linear shift, we only have to adjust the first-order derivative $\partial f/\partial x$ as follows:
$\frac{\partial\tilde{f}_{\text{shift}}(x)}{\partial x}=\frac{\partial\tilde{f}(x)}{\partial x}-m,$ (49)
where $m$ is the angular coefficient associated with $\text{shift}(x)$.
B.3 Simple example in the multivariate case
Let |$\ket{\psi_{g}}\in(\mathbb{C}^{2})^{\otimes 6}$ be a six-qubit state and let $\lambda_{g}=7$ be a scalar used to model a bivariate solution function $g:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, such that :
|$\ket{\psi_{g}}=\frac{1}{\sqrt{14}}\left(2\ket{001000}+2\ket{100010}+\sqrt{6}\ket{010111}\right)~{},$
with
$g(x,y)=\lambda_{g}\langle\psi_{g}|O_{C}(x,y)|\psi_{g}\rangle$
We choose to encode the $x$-dependence with three qubits and the $y$-dependence with two qubits. Therefore, we can divide any basis state |$\ket{i}$ of six-qubit systems as follows:
[…]
with $\textrm{bin}(L_{1})=i_{1}i_{2}i_{3}$ and $\textrm{bin}(L_{2})=i_{4}i_{5}$. If we take for instance the first basis state appearing in |$\ket{\psi_{g}}$, we can associate it with the following multivariate Chebyshev polynomial after computation of the expectation value of the diagonal observable $O_{C}(x,y)$:
|$\underbrace{\ket{0}}_{+}\underbrace{\ket{010}}_{2}\underbrace{\ket{00}}_{0}~{}\xrightarrow{\langle O_{C}(x,y)\rangle}~{}\text{Cheb}(2,x)\text{Cheb}(0,y)=2x^{2}-1~{},$
with
$O_{C}(x,y)=\text{diag}\Big{(}\text{Cheb}(0,x,y),\text{Cheb}(1,x,y),\dots,-\text{Cheb}(30,x,y),-\text{Cheb}(31,x,y)\Big{)}~{},$
which is equal to:
$\begin{split}O_{C}(x,y)=\text{diag}\Big{(}\text{Cheb}(0,x)\text{Cheb}(0,y),\text{Cheb}(0,x)\text{Cheb}(1,y),\dots,\\ -\text{Cheb}(7,x)\text{Cheb}(2,y),-\text{Cheb}(7,x)\text{Cheb}(3,y)\Big{)}~{}.\end{split}$
Therefore, we can express the solution function $g$ as follows:
$g(x,y)=\frac{7}{14}\Big{(}4\cdot\text{Cheb}(2,x)\text{Cheb}(0,y)-4\cdot\text{Cheb}(0,x)\text{Cheb}(2,y)+6\cdot\text{Cheb}(5,x)\text{Cheb}(3,y)\Big{)}~{}.$
We can rewrite the univariate polynomials to obtain:
$g(x,y)=\frac{1}{2}\Big{(}4(2x^{2}-1)-4(2y^{2}-1)+6(16x^{5}-20x^{3}+5x)(4y^{3}-3y)\Big{)}~{}.$
After some simplifications, we conclude that:
$g(x,y)=4x^{2}-4y^{2}-45xy+180x^{3}y+60xy^{3}-144x^{5}y-240x^{3}y^{3}+192x^{5}y^{3}~{}.$ (50)
Appendix C Run time analysis details
This section details the time taken to run various parts of the algorithm.
C.1 Generating the observables
In this subsection, we study the function called in line 11 of Alg. 1.
Algorithm 2 Observables generation function
1:$G_{F}$ : the set of orders of derivatives involved in $E$ for each function in $F$
2:$D$ : the set of domains of definition for each function in $F$
3:$n$ : the number of qubits
4:$n_{s}$ : the number of samples
5:the list of observables capable of computing the value of any function’s derivative at any sample point, where the function is encoded in the output state of our VQA
6:function GenerateObservables($G_{F},D,n,n_{s}$)
7:     $G\leftarrow\cup_{f\in F}G_{F}(f)$
8:     $S\leftarrow$ generateSamples($n_{s}$, $D$)
9:     listObs $\leftarrow$ [ ]
10:     for $g$ in $G$ do
11:         $\text{order}_{g}$leftarrow [ ]
12:         for $x_{s}$ in $S$ do
13:              C $\leftarrow$ [ ]
14:              for i from 0 to $12^{n-1}-$ 1 do
15:                  C.append(Cheb($i$, $x_{s}$, $g$))
16:              end for
17:              observable $\leftarrow$ diag(C, --C)
18:              $\text{order}_{g}$.append(observable)
19:         end for
20:         listObs.append($\text{order}_{g}$)
21:     end for
22:     return listObs, $S$
23:end function
$T_{obs}=T_{union}+T_{samples}+T_{der\_iters}~{}.$
In Alg. 2:
  • In line 7, we deduce all the derivation levels in the DEs in time $T_{union}$. This step is just a union of sets, so we assume that $T_{union}\ll T_{obs}$.
  • In line 8, we compute all the sample points used for the evaluation in $T_{samples}$. This step consists of the creation of a list of regularly spaced points, and we assume that $T_{union}\ll T_{samples}$.
  • In line 10, we iterate over each order of derivation (0, 1, …) in $T_{der\_iters}=|G|\times T_{samp\_iter}$.
  • In line 12, we iterate over each sample point in $T_{samp\_iters}=n_{s}\times(T_{order\_iter}+T_{diag})$.
  • In line 14, we iterate over each Chebyshev polynomial’s order and compute the corresponding value for the Chebyshev monomial in $T_{order\_iter}=2^{n-1}\times T_{cheb}$. We can remark here that it is possible to enhance this part by not computing the evaluation of Chebyshev polynomials that are used for several orders twice.
  • In line 15, we compute the value of the Chebyshev monomial for a given order, sample value, and derivative order in $T_{cheb}$.
  • In line 17, we instantiate the diagonal observable, with the elements on the diagonal being the input of the function. This can be done by instantiating a matrix or by using the Pauli string decomposition of the observable given in Eq. C.1. This step takes a time $T_{diag}$ (which is linearly dependent on the size of the input).
The overall time is then:
$T_{obs}=|G|\times n_{s}\times(2^{n-1}\times T_{cheb}+T_{diag})$ (51)
Some theory: The observable in this algorithm can be written as:
$O_{C}(x)=\sigma_{z}\otimes\left(\sum_{i=0}^{2^{n-1}-1}\text{Cheb}(i,x)\outerproduct{i}{i}\right)$
$=\begin{pmatrix}\text{Cheb}(0,x)&&&&&\\ &\ddots&&&&\text{\huge 0}\\ &&\text{Cheb}(2^{n-1}-1,x)&&&\\ &&&-\text{Cheb}(0,x)&&\\ \text{\huge 0}&&&&\ddots&\\ &&&&&-\text{Cheb}(2^{n-1}-1,x)\\ \end{pmatrix}~{}.$ (52)
In order to generate the full matrix, we construct the diagonal of the matrix using the tensor product of linear combinations of the Pauli matrix $Z$ and the identity $I=I_{2}$, as well as ${\rm bin}(i)$, the binary representation of $i$:
|$\outerproduct{i}{i}=\bigotimes_{k\in{\rm bin}(i)}\frac{1}{2}(I+(\shortminus 1)^{k}Z)~{}.$ (53)
Then, using the tensor products, it takes $n(T_{tensor,2\times 2}+T_{sum,2\times 2})$ to construct $O_{C}(x)$ (per given point $x$), with $T_{tensor,2\times 2}$ the time required to compute the tensor product of a matrix with a $2\times 2$ matrix, and $T_{sum,2\times 2}$ the time to add or subtract two $2\times 2$ matrices.
C.2 Variational quantum circuit (VQC)
We use the so-called hardware-efficient Ansatz [CAB+21] to construct the state encoding the solution function. Each layer contains parameterized single-qubit rotations for all the qubits, which are then entangled by CNOT gates in a braided (or alternating) fashion.
[…]
Figure 6: An example of our VQC with $n=5$ qubits and depth $d=2$.
There are two types of times attached to the VQC: the initialization of the circuit and the execution of the circuit. The initialization takes a time
$T_{VQC\_init}=(2n-1)\times d\times(T_{gate\_creation}+T_{instruction\_adding})$ (54)
and the execution of the circuit takes a time
$T_{VQC\_exec}=3\times d\times T_{gate\_exec}.$ (55)
One can remark that, even in the presence of several functions in the differential equations, all the circuits can be run in parallel. When the number of qubits and the depth are the same for all functions, we can also imagine only using a single circuit that is initialized once and just replacing the angles to evaluate one function or another.
C.3 Expectation values
In order to determine the value of the solution $f$ at a given point $x$, we compute the expectation value with respect to the diagonal observable $O_{C}(x)$ discussed above as follows:
$𝜓𝜃f_{\theta}(x)=\lambda~{}\expectationvalue{O_{C}(x)}{\psi(\theta)}.$ (56)
Though one has to deal only with diagonal matrices, evaluating the expectation values of several observables constitutes the costly step of the quantum part of the algorithm. The method of expressing the elements of the matrices is convenient since evaluating the expectation values of Pauli strings through a quantum computer is practical, especially if the Pauli strings commute, since in that case they can be measured simultaneously [KRP22]. The upper bound of the number of such strings per observable is $2^{n-1}$, with $n$ the number of qubits. So in the worst case scenario, in which no strings commute, one has $T_{eval}=2^{n-1}T_{measurement}$.
But in our case, our Pauli strings have a very specific form: they all start with a Pauli $Z$ gate and can then take all combinations of Pauli $I$ and $Z$. In that case, all the Pauli operators used to decompose our diagonal observables commute. Therefore, one can imagine many protocols and methods for efficiently evaluating the expectation values of such observables, even in the presence of noise [ZNCA, KSK+22, HKP21, CG23a, HBRM20, KIK+22, RB20, HKP20, KG22]. Once we know all the expectation values of the Pauli observables, we can reconstruct all expectation values of all diagonal observables (for all samples, for evaluating solution functions or their derivatives) in parallel with classical post-processing.
The evaluation of the expectation values consists of running the circuit and carrying out the right measurements for each shot, then post-processing them to reconstruct all desired expectation values, giving
$T_{expvalue}=n_{shots}\times(T_{VQC\_exec}+T_{measurement})+T_{post\_estim}.$ (57)
C.4 Loss function (error)
The value we want to minimize is:
$L^{\rm{diff}}(\theta)=\frac{1}{n_{s}}\sum_{e\in E}\sum_{x_{s}\in S}e(x_{s})^{2}~{},$ (58)
where the notation is the same as that defined in Sec. 2.5.
We omit the term in the loss that concerns the boundary conditions and apply floating boundary conditions handling. Recall that we have:
  • $\theta$: the set of angles parametrizing our solution,
  • $E$: the set of differential equations,
  • $n_{s}$: the number of samples,
  • $F$: the set of functions,
  • $n_{f}$: the number of functions involved in the equations,
To compute this sum, we iterate over:
  1. 1.
    the set of differential equations $E$,
  2. 2.
    the set of samples ($n_{s}$ elements),
  3. 3.
    the set $H_{E}$ of all solution functions and their derivatives appearing in the differential equations.
Given this, the time to compute this loss value in the worst case (for all functions, with the function and its derivatives appearing in each equation) would be:
$T_{loss}=|E|\times n_{s}\times(T_{square}+\sum_{f\in F}|H_{E}|T_{sum}),$ (59)
where $T_{square}$ is the time required to square a floating point number and $T_{sum}$ is the time needed to sum two floating point numbers.
Appendix D Performance profile
Performance profiles, as described in [DM02], are a fairly common tool in optimization when it comes to comparing the performance of different algorithms, software, or methods for solving a problem. They allow for assessing both the speed and accuracy of an algorithm in relation to the proposed solutions. They can also be used to compare the performance of a given solver with different parameters and settings.
We denote the set of solvers to compare as $\mathcal{S}$ and the set of problems on which the solvers will be compared as $\mathcal{P}$. In the context of this paper, the problems to solve correspond to different sets of partial DEs. One then usually needs to choose a criterion to measure the performance or quality of the solver. It can be the number of evaluations of the solution function, the final value of the loss function, the precision of the result, or the number of iterations needed for a given precision. In our case, the validation score we defined previously is used as the measure of performance for a given solver.
We thus define the quantity $V_{p,sol}$, which is the validation score of the solver $sol\in\mathcal{S}$ for solving the problem $p\in\mathcal{P}$. In order to be able to compare several solvers, we need to have a common basis for comparison. As detailed in [DM02], for a given problem $p$, we compare the performance of each solver with the best performance over $\mathcal{S}$. We then define the performance ratio $r_{p,sol}$ as follows:
$r_{p,sol}=\frac{V_{p,sol}}{\min\limits_{sol\in\mathcal{S}}V_{p,sol}}~{}.$ (60)
Therefore, for a given problem $p$, we define the performance ratio as the validation score of the solver for this particular problem divided by the best validation score (the minimum) among all other solvers, still for the same problem. We determine that the solver could not solve the problem if the validation score is above the acceptance threshold. If a solver is not able to solve a problem, the performance ratio is set to a maximum value $r_{M}$.
We can now define the global performance of a solver $sol$ for all problems in $\mathcal{P}$ as the cumulative distribution function
$\rho_{sol}(\tau)=\frac{\text{card}\Big{(}\{p\in\mathcal{P}~{}|~{}r_{p,sol}\leq\tau\}\Big{)}}{\text{card}(\mathcal{P})}~{},$ (61)
where $\rho_{sol}(\tau)$ is the probability for the solver $sol$ that a performance ratio $r_{p,sol}$ is within a factor $\tau\in\mathbb{R}$ of the best possible ratio (which is 1). The function $\rho_{sol}$ can be seen as the (cumulative) distribution function for the performance ratio. To obtain the performance graph of the solver $sol$, we will represent the strictly increasing function $\rho_{sol}:\mathbb{R}\to[0,1]$ as a function of $\tau$.
The value $\rho_{sol}(1)$ is the probability that the solver $sol$ has a better validation score than all the other solvers. Therefore, if we are interested in precision, we can only compare the values of $\rho_{sol}(1)$ for all solvers in $\mathcal{S}$.
Furthermore, we recall that any performance ratio takes a value within $[1,r_{M}]$, and that $r_{p,sol}=r_{M}$ only if the problem $p$ was not solved by $sol$. We can deduce that $\rho_{sol}(r_{M})=1$ and that the quantity
$\rho_{sol}^{*}=\lim_{\tau\rightarrow{r_{M}}^{-}}\rho_{sol}(\tau)$ (62)
is the probability that $sol$ is able to solve a problem. Thus, if we are only interested in solvers that have the highest probability of success, we compare values of $\rho_{sol}^{*}$ for all solvers in $\mathcal{S}$ and keep that with the highest value. The value of $\rho_{sol}^{*}$ can be easily observed in a performance profile graph because the value of $\rho_{sol}$ stagnates for high values of $\tau$.
Figure 7: Example of a performance profile graph from [DM02]. The highest value of $\rho_{sol}(1)$ is here achieved by LOQO (0.61), while the highest value of $\rho_{sol}^{*}$ is apparently achieved by MINOS (0.8) for $\tau\in[1,10]$.